Wellstone Inc. produces and markets replacement covers for cell phones in a variety of colors. The company would like to allocate its production plans to five different colors: bright white, metallic black, magnetic lime, tangerine orange, and fusion red. The company set up a kiosk in the Mall of America for several hours and asked randomly selected people which cover color was their favorite. The results follow:

   Bright white Metallic black Magnetic lime Tangerine orange Fusion red

130 104 325 455 28

 Draw a bar chart & pie chart for the table.  

If Wellstone Inc. plans to produce 12,000 cell phone covers, how many of each color should it produce?



1
Expert's answer
2021-02-15T17:50:56-0500

A bar chart will show the bars with height of 130, 104, 325, 455 and 28, and a pie chart will show the persentage parts of every color.

130 + 104 + 325 + 455 + 28 = 1042,

12,000/1042 = 11.52, so if Wellstone Inc. plans to produce 12,000 cell phone covers, the it will produce:

130×11.52 = 1497 of bright white,

104×11.52 = 1198 of metallic black,

325×11.52 = 3743 of magnetic lime,

455×11.52 = 5240 of tangerine orange,

28×11.52 = 322 of fusion red.